A ticket booth sold 226 tickets and collected $843 in ticket sales. Adult tickets are $5.50 and child tickets are $1.50. How many tickets of each type were sold?

Found 2 solutions by rm29924, checkley79:
Answer by rm29924(97)

Let A=the amount of adult tickets sold
Let C=the amount of child tickets sold
their sum should be 226
A+C=226
5.50A+1.50C=843
solve for A in the first equation
A=226-C
plug that value for A into the second equation
5.50(226-C)+1.50C=843
1243-5.50C+1.50C=843
1243-4C=843
400=4C
C=100
plug that into the first equation
A+C=226
A+100=226
A=126
there were 100 child tickets and 126 adult tickets sold
you can check your work by plugging the answers into the second equation



Answer by checkley79(3341)

A+C=226
A=226-C
5.5A+1.50C=843
5.5(226-C)+1.50C=843
1143-5.5C+1.50C=843
-4C=843-1143
-4C=-400
C=-400/-4
C=100 CHILDREN'S TICKETS WERE SOLD.
A+100=226
A=226-100
A=126 ADULT TICKETS SOLD.
PROOF:
5.5*126+1.50*100=843
693+150=843
843=843